OpenStudy (anonymous):

How many atoms are in 65 grams of argon?

OpenStudy (anonymous):

n= m/M n=N/L where: m - mass M - molar mass N - number of atoms/molecules L - avogadro constant so: N= m*L/M

OpenStudy (anonymous):

can you help me through it please?

OpenStudy (anonymous):

well i have written it all now all you need to do is to put in numbers

OpenStudy (anonymous):

i honestly have no clue how to even start the problem

OpenStudy (anonymous):

N= 65 g * 6,022 * 10^23 mol-1 / 39,9 g mol-1

OpenStudy (anonymous):

is it 9.79x10^23?

OpenStudy (anonymous):

yes i think it is...